OpenStudy (xxdragoonxx):

How much weather stripping wall will you need to go around a window that is 6 7/8 feet long and 2/14 feet wide?

OpenStudy (sepeario):

the formula for the perimeter of this figure is 2(l+w), where l and w are length and width respectively. hence: \[2(6\frac{ 7 }{ 8 }+\frac{ 2 }{ 14 }) = 2(7\frac{ 1 }{ 56 })\]

OpenStudy (sepeario):

So the answer is 14 2/56 or 14 1/28.